OpenStudy (anonymous):

Solve by substitution: 2x+y=0 x-3y=-7 please help.

OpenStudy (anonymous):

Please help?

OpenStudy (anonymous):

substitution requires you to solve for one variable in one equation, and substitute your results into the other equation. for example, if you solve the first equation for 'y' you'll obtain: y = -2x. then you replaced 'y' with '-2x' in the other equation and solve for x.

OpenStudy (anonymous):

replace*

OpenStudy (anonymous):

So How do I break it down to actually show my math? This is where I get stuck Lol. I understand what you're saying. It's just showing my work is where I get in trouble.

OpenStudy (anonymous):

2x+y=0....1) x-3y=-7....2) y=-2x...3) therefore: substitute 3) into 2) x-3(-2x)=-7 x+6x=-7 7x=-7 x=-1 therefore:y=-2x y=-2(-1) y=2

OpenStudy (anonymous):

So y =2 got that. This makes sense now. Thank you so much.